The aluminum wire has a cross-sectional area of 40 mm
2
. The voltage at its ends is 3 mV.
Какова длина провода, если по нему течет ток силы 0,5 А? What is the length of the wire if a current
of 0.5 A flows through it? The electrical resistivity of aluminum is 0,026 ·m.
1
Expert's answer
2020-11-23T10:28:23-0500

The resistance of wire is

"R=\frac{U} {I} =\rho\frac{l} {S},"

where U - the voltage at the ends of wire,

I - the current through the wire,

"\rho" - resistivity of aluminium,

"l" - length of the wire,

S - cross-sectional area of the wire.

Length of the wire from equation above is

"l=\frac{US} {\rho I} = \frac {3\cdot 10^{-3}\cdot 40 \cdot 10^{-6} }{0.026\cdot 0.5}=9.23\cdot 10^{-6} \space m"

Answer: "9.23\cdot 10^{-6} \space m"